# Audio Super-Resolution with Latent Bridge Models

### v2

@@ A.2 Importance of $f_{\text{prior}}$ detection for inference
At inference time, $f_{\text{prior}}(x)$ is estimated for each input sample with the same detection method of $f_{\text{eff}}$ individually, allowing adaptive filtering tailored to the true spectral bandwidth of the audio. This ensures that super-resolution is performed on the informative signal subspace, leading to improved perceptual quality and generalization When generative models are used as the source of input audio, such as MaskGCT [111], the nominal output sampling rate is fixed (e.g., 24 kHz). However, the actual frequency content may vary depending on the speech prompt or training data, and may not occupy the full bandwidth of the input sampling rate. Similar phenomena have also been observed in 16 kHz generation models [61], where the effective bandwidth is often narrower than the nominal sampling rate. Consequently, using the declared sampling rate as the low-pass cutoff for super-resolution input may lead to incorrect reconstruction.
Figure 9: Super-resolution results for samples generated by MaskGCT [111] at 24 kHz, using both AudioSR and our proposed system. While the nominal sampling rate is 24 kHz, the actual spectral content does not span the full bandwidth. As shown in the result of AudioSR, using 24 kHz directly as the cutoff frequency results in incorrect high-frequency generation.
@@ Appendix B Compression network
Variational autoencoders (VAEs) [42] include an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder maps an input $\bm{x}$ to a latent representation $\bm{z}=\mathcal{E}(\bm{x})$, and the decoder reconstructs $\bm{x}$ from $\bm{z}$ via $\hat{\bm{x}}=\mathcal{D}(\bm{z})$. The training loss is
@@
We follow the similar training network structure and pipeline with Stable-Audio-Open [28] and ETTA [50], which combines the following training objectives:
@@
    A multi-resolution STFT loss [99, 100]:
@@
    An adversarial hinge loss and feature matching loss from EnCodec [21]:
@@ B.1.1 Compression rate
As shown in Table 4, for the generation task, higher compression typically improves generative modeling without significantly compromising reconstruction [26]. However, in low-level tasks such as audio super-resolution, we observe that lower compression rates consistently yield better reconstruction and SR performance. Nevertheless, lower compression introduces longer training time and higher inference cost (RTF), thus reducing deployment efficiency.
@@ B.1.2 KL divergence
For diffusion-based generation, increasing the KL divergence term is known to improve the alignment between the latent space and the Gaussian prior, thereby enhancing the diffusability of autoencoders and improving generation quality [117, 97]. However, in bridge models, the prior distribution is no longer a standard Gaussian but instead a Dirac distribution induced by the low-resolution waveform latent. In this setting, enforcing KL regularization may no longer be appropriate. Moreover, a smaller—or even zero—KL weight allows the latent space to prioritize reconstruction fidelity, potentially lifting the upper bound of super-resolution quality.
@@ B.1.3 Comparison with baselines
100 Hz agc [3] (continuous) 0.913 0.762 0.766 0.741
50 Hz agc [3] (discrete) 0.917 0.789 0.754 0.768
150 Hz encodec [21] 0.887 1.126 0.761 1.176
75 Hz audiodec [115] 0.922 0.939 0.944 0.937
75 Hz flowdec [114] 0.925 0.619 0.560 0.630
@@
100 Hz agc [3] (continuous) 0.730 0.836 0.748 0.840
50 Hz agc [3] (discrete) 0.741 0.852 0.703 0.862
150 Hz encodec [21] 0.704 0.925 0.910 0.922
100 Hz dac [48] 0.734 0.799 0.735 0.801
75 Hz audiodec [115] 0.695 0.999 0.969 0.997
75 Hz flowdec [114] 0.717 0.899 0.810 0.908
@@ Appendix C Foundation of bridge models
Bridge models provide a principled and efficient solution for inverse problems and conditional generation in settings where the prior is non-Gaussian yet partially aligned with the target distribution. Given a sample from the prior distribution $x_{\text{prior}}$ and its correspondence from target distribution $x_{\text{target}}$, In score-based generative models (SGMs) [98], a forward SDE is defined between $x_{0}=x_{\text{target}}\sim p_{\text{target}}$, and $x_{T}=x_{\text{prior}}\sim p_{\text{prior}}$:
@@
Under the above framework with boundary constraints, the Schrödinger bridge (SB) problem [94, 14] seeks to find a path measure $p$ of specified boundary distributions that minimizes the Kullback-Leibler divergence $D_{\text{KL}}$ between a path measure and reference path measure $p_{\text{ref}}$:
@@
In SB theory [108, 13, 14], this specific SB problem can be expressed as a pair of forward-backwords linear SDEs:
@@
where the non-linear drifts $\nabla\log\Psi_{t}(x_{t})$ and $\nabla\log\widehat{\Psi}_{t}(x_{t})$ can be described by coupled partial differential equations (PDEs). A closed-form solution for SB [14] exists when Gaussian smoothing is applied with $p_{0}=\mathcal{N}\left(x_{\text{HR}},\epsilon_{0}^{2}I\right)$ and $p_{T}=\mathcal{N}\left(x_{\text{LR}},\epsilon_{T}^{2}I\right)$, to the original Dirac distribution. By defining $\alpha_{t}=e^{\int_{0}^{t}f(\tau)d\tau}$, $\bar{\alpha}_{t}=e^{\int_{1}^{t}f(\tau)d\tau}$, $\sigma_{t}^{2}=\int_{0}^{t}\frac{g^{2}(\tau)}{\alpha_{\tau}^{2}}d\tau$, and $\bar{\sigma}_{t}^{2}=\int_{t}^{1}\frac{g^{2}(\tau)}{\alpha_{\tau}^{2}}d\tau$, with boundary conditions and linear Gaussian assumption, tractable form of SB is solved as
@@
under $\epsilon_{T}=e^{\int_{0}^{T}f(\tau)d\tau}\epsilon_{0}$ and $\epsilon_{0}\rightarrow 0$ [14]. Therefore, marginal distribution of $x_{t}$ at state $t$ is also Gaussian:
@@ D.1 Waveform-based filtering augmentation
Figure 10: Visual comparison of waveform-based filtering strategies. (a) shows the baseline output from the any-to-48 kHz stage with high-frequency artifacts. (b) demonstrates that direct super-resolution from such inputs results in degraded high-frequency reconstruction. (c)–(f) apply low-pass filters with different cutoff frequencies, and the corresponding outputs after super-resolution are shown in (g)–(i).
Samples generated from the any-to-48 kHz stage often exhibit small-scale high-frequency artifacts, as shown in Figure 10(a). Directly feeding such artifacts into the next-stage super-resolution model can significantly degrade performance, as shown in Figure 10(b), where high-frequency content becomes difficult to reconstruct. To address this, we apply Chebyshev Type-I low-pass filters (order=8) with sharp frequency roll-off in the spectral domain, as shown in Figure 10(c)–(f), removing a frequency content of 8 kHz, 4 kHz, and 2 kHz, respectively, before feeding them into the next-stage model. The resulting outputs are shown in Figure 10(g)–(i).
@@ E.2 Low-frequency replacement
Low-frequency replacement is a widely adopted post-processing technique in super-resolution systems, particularly in super-resolution models based on compressed representations [60, 62, 122, 37].
@@ E.3 Inference computational cost
Bridge-SR [54] 48 kHz Waveform Bridge 1.670 50
UDM+ [60] 48 kHz Waveform Unconditional Diffusion 2.320 100
AudioSR [62] 48 kHz Mel-VAE Conditional Diffusion 0.948 50
Fre-painter [41] 48 kHz Mel GAN 0.009 1
NVSR [60] 48 kHz Mel GAN 0.033 1
@@
It is worth noting that under the 48 kHz setting, although our method is not as fast as GAN-based approaches such as [60, 41], due to the iterative sampling nature, it significantly outperforms other iterative-based methods including [54] and the previous state-of-the-art system AudioSR [62].
When scaling to higher sampling rates (e.g., 192 kHz), it naturally leads to slower inference speeds. We mitigate this impact by adopting a lighter network architecture (see Appendix G.2), so the inference speed does not increase linearly with the sampling rate. As shown in Table 12, even when upsampling to 192 kHz, our system is still faster than the 48 kHz waveform-domain bridge system [54].
@@ E.4 Modeling space and other probabilistic methods
To isolate the contribution of the latent-bridge formulation, we conducted additional experiments using the same speech datasets (OpenSLR [43], Expresso [75], EARS [86], and VCTK-Train [118]) for training all models, and evaluated them on the VCTK-Test set. We compare the following variants without any additional modules, each trained with 500K steps on the any-to-48 kHz setting and tested on the 8-to-48 kHz SR setting with 50-step sampling:
@@
  • Bridge-STFT (based on the Nemo [33] framework)
  • Bridge-Waveform (based on Bridge-SR [54])
@@
Method Modeling Space SSIM $\uparrow$ LSD $\downarrow$ LSD-LF $\downarrow$ LSD-HF $\downarrow$ SigMOS [87] $\uparrow$
@@ G.1 Training dataset
48 kHz VCTK-train [118] 40h Speech
OpenSLR [43] 190h Speech
EARS [86] 100h Speech
Expresso [75] 20h Speech
MusDB18 [82] 10h Music
Medleydb [6] 10h Music
@@ G.2 Model architecture
We adopt the Diffusion Transformer (DiT) architecture as the noise predictor for our bridge model, following the design of Stable Audio Open [28]. During training, we employ the Bridge-gmax schedule, which has been proven effective in previous super-resolution and text-to-speech tasks [54, 14]. The network architecture and training configurations for the three-stage cascade are detailed below.
@@ G.3 Evaluation metrics
While LSD is a point-wise spectral metric, Structural Similarity (SSIM) [113] overcomes its limitations by incorporating structural and contextual information. Originally designed for perceptual image quality assessment, SSIM compares local statistics of luminance, contrast, and structure. In the context of audio, we apply SSIM on log-magnitude spectrograms to capture spectral texture fidelity.
@@
Here, $S_{k}$ and $\hat{S}_{k}$ denote the $k$-th local $7\times 7$ block in the spectrograms, $K$ is the total number of blocks, and $\epsilon_{1}=0.01$, $\epsilon_{2}=0.02$ are stability constants following VoiceFixer [24]. Higher SSIM values imply better preservation of spectral structure.
@@
To assess perceptual speech quality in an objective manner, we adopt the Virtual Speech Quality Objective Listener (ViSQOL) [15], a signal-based estimator of Mean Opinion Score (MOS) using spectro-temporal similarity.
ViSQOL (in audio mode) assumes a sampling rate of 48 kHz and outputs MOS-LQO scores ranging from 1 to 4.75, with higher scores indicating better perceptual quality. However, since ViSQOL is designed for speech and only supports certain sample rates (e.g., 16 kHz), we additionally employ SigMOS [87] to evaluate speech quality for 48 kHz scenario.
@@ G.4 Subjective experiment
For the subjective listening test, we evaluate super-resolution performance on generative model outputs from three representative systems: MaskGCT [111], AudioLDM2 [63], and QA-MDT [53], which represent modern speech generation at 24 kHz and music/sound generation at 16 kHz, respectively.
For speech, we randomly select 10 samples from the LibriSpeech-test set [77] as both the reference audio and the text prompt for MaskGCT generation. For music and environmental audio, we sample 10 examples each from the MusicCaps [1] and AudioCaps [40] datasets. To further assess performance on real-world music, we additionally include 10 music segments from the Song-Describer dataset [68]. All audio clips are truncated to 5.12 seconds. For each test case, we present three versions to listeners: the input audio (Generated), the output from AudioSR, and the output from our system (AudioLBM), yielding a total of 40 groups × 3 samples = 120 audio clips for evaluation.
@@ H.1 Detailed introduction of audio super-resolution baselines
Nu-Wave [49] represents the first attempt at applying diffusion models to waveform-level audio super-resolution. It builds on the WaveNet [106] structure and adopts the DiffWave [44] training paradigm to learn mappings from fixed low-resolution audio to 48 kHz waveform. To support evaluation across various input resolutions, we follow the official implementation [^3] https://github.com/maum-ai/nuwave and train separate models for 8 kHz, 12 kHz, 16 kHz, and 24 kHz inputs, each for 1.5 million steps.
@@
Nu-Wave 2 [49] extends Nu-Wave by incorporating two key innovations: Short-Time Fourier Convolution (STFC) and Bandwidth Spectral Feature Transform (BSFT). These additions improve harmonic modeling and allow support for any-to-48 kHz super-resolution. The model also adopts a fast generation scheme using eight predefined non-uniform sampling steps. For fair comparison, we use the pre-trained model made available by the authors [^4] https://github.com/maum-ai/nuwave2.
@@
UDM+ [120] retains the DiffWave-based architecture but deviates from Nu-Wave in training strategy. It uses unconditional diffusion modeling and injects low-resolution signal guidance at inference through a 50-step uniform sampling process with low-frequency component replacement[67]. Additionally, Manifold Constraint Gradient (MCG) [17] is introduced to better balance low- and high-frequency consistency during generation.
@@
Bridge-SR [54] adapts the Nu-Wave 2 architecture but introduces a key conceptual shift: instead of sampling from a standard Gaussian prior, the model initializes directly from the low-resolution input. This bridge formulation allows high-quality any-to-48 kHz SR. Furthermore, it employs data normalization strategies and frequency-aware loss functions to improve the final reconstruction quality.
@@
mdctGAN [96] targets the instability issues of complex-valued neural networks in the STFT domain by operating in the Modified Discrete Cosine Transform (MDCT) domain. Rather than relying on post-vocoders, mdctGAN uses adversarial learning to generate high-fidelity waveform with phase consistency. We evaluate the model using the authors’ official checkpoints [^5] https://github.com/neoncloud/mdctGAN, covering four different input resolutions.
@@
NVSR [60] is a neural vocoder-based speech SR system tailored to handle a wide range of upsampling factors and phase reconstruction challenges. It is composed of three main modules: (1) a mel-bandwidth extension network based on ResUNet; (2) a TFGAN-powered [105] neural vocoder; and (3) a post-processing module. We use the official model checkpoint available at  [^6] https://github.com/haoheliu/ssr_eval for our comparisons.
@@
AP-BWE [65] is the first bandwidth extension system to model the high-frequency phase explicitly. It leverages a GAN-based framework to jointly predict amplitude and phase spectra using a dual-stream CNN, where the two branches interact to reconstruct full-band audio from narrowband speech. Although it does not support arbitrary-resolution inputs, we evaluate the four fixed-resolution models provided by the authors [^7] https://github.com/yxlu-0102/AP-BWE.
@@
AudioSR [62] is a general-purpose diffusion-based super-resolution model capable of handling speech, music, and general audio inputs. It performs super-resolution in the latent space conditioned on low-resolution inputs and uses a two-stage cascade—consisting of a mel-spectrogram VAE and a vocoder—for waveform reconstruction. Benefiting from strong zero-shot generalization, we evaluate the model using the official implementation [^8] https://github.com/haoheliu/versatile_audio_super_resolution.
@@
Fre-Painter [41] introduces a masked autoencoder (MAE)-based framework for audio super-resolution. It employs an upper-band masking strategy during fine-tuning to simulate low-resolution inputs by masking high-frequency components. Additionally, a mix-ratio masking approach is utilized to enhance robustness across various input sampling rates. The model is pre-trained on large-scale datasets to learn robust speech representations and is fine-tuned jointly with a neural vocoder for waveform reconstruction. We evaluate the model using the official implementation [^9] https://github.com/FrePainter/code.
@@
FLowHigh [122] proposes an audio super-resolution method based on flow matching, aiming to overcome the sampling inefficiency of traditional diffusion models. It adopts an inpainting-style framework similar to AudioSR but constrains the modeling space to Mel-spectrograms. A lightweight two-layer Transformer [107] is used as the velocity predictor. To further improve sampling efficiency, FLowHigh introduces a data-dependent prior instead of a pure Gaussian initialization, allowing for faster flow-based sampling. We evaluate the model using the official implementation [^10] https://github.com/jjunak-yun/FLowHigh_code.

### v3

@@ A.2 Importance of $f_{\text{prior}}$ detection for inference
At inference time, $f_{\text{prior}}(x)$ is estimated for each input sample with the same detection method of $f_{\text{eff}}$ individually, allowing adaptive filtering tailored to the true spectral bandwidth of the audio. This ensures that super-resolution is performed on the informative signal subspace, leading to improved perceptual quality and generalization When generative models are used as the source of input audio, such as MaskGCT [wang2024maskgct], the nominal output sampling rate is fixed (e.g., 24 kHz). However, the actual frequency content may vary depending on the speech prompt or training data, and may not occupy the full bandwidth of the input sampling rate. Similar phenomena have also been observed in 16 kHz generation models [liu2023audioldm], where the effective bandwidth is often narrower than the nominal sampling rate. Consequently, using the declared sampling rate as the low-pass cutoff for super-resolution input may lead to incorrect reconstruction.
Figure 9: Super-resolution results for samples generated by MaskGCT [wang2024maskgct] at 24 kHz, using both AudioSR and our proposed system. While the nominal sampling rate is 24 kHz, the actual spectral content does not span the full bandwidth. As shown in the result of AudioSR, using 24 kHz directly as the cutoff frequency results in incorrect high-frequency generation.
@@ Appendix B Compression network
Variational autoencoders (VAEs) [kingma2013auto] include an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder maps an input $\bm{x}$ to a latent representation $\bm{z}=\mathcal{E}(\bm{x})$, and the decoder reconstructs $\bm{x}$ from $\bm{z}$ via $\hat{\bm{x}}=\mathcal{D}(\bm{z})$. The training loss is
@@
We follow the similar training network structure and pipeline with Stable-Audio-Open [evans2025stable] and ETTA [lee2024etta], which combines the following training objectives:
@@
    A multi-resolution STFT loss [steinmetz2020auraloss, steinmetz2021automatic]:
@@
    An adversarial hinge loss and feature matching loss from EnCodec [defossez2022high]:
@@ B.1.1 Compression rate
Typically, for the generation task, higher compression typically improves generative modeling without significantly compromising reconstruction [evanslong]. However, in low-level tasks such as audio super-resolution, we observe that lower compression rates consistently yield better reconstruction and SR performance, as shown in Table 4. Nevertheless, lower compression introduces longer training time and higher inference cost (RTF), thus reducing deployment efficiency.
@@ B.1.2 KL divergence
For diffusion-based generation, increasing the KL divergence term is known to improve the alignment between the latent space and the Gaussian prior, thereby enhancing the diffusability of autoencoders and improving generation quality [xu2025exploring, skorokhodov2025improving]. However, in bridge models, the prior distribution is no longer a standard Gaussian but instead a Dirac distribution induced by the low-resolution waveform latent. In this setting, enforcing KL regularization may no longer be appropriate. Moreover, a smaller—or even zero—KL weight allows the latent space to prioritize reconstruction fidelity, potentially lifting the upper bound of super-resolution quality.
@@ B.1.3 Comparison with baselines
100 Hz agc [agc2024] (continuous) 0.913 0.762 0.766 0.741
50 Hz agc [agc2024] (discrete) 0.917 0.789 0.754 0.768
150 Hz encodec [defossez2022high] 0.887 1.126 0.761 1.176
75 Hz audiodec [wu2023audiodec] 0.922 0.939 0.944 0.937
75 Hz flowdec [welker2025flowdec] 0.925 0.619 0.560 0.630
@@
100 Hz agc [agc2024] (continuous) 0.730 0.836 0.748 0.840
50 Hz agc [agc2024] (discrete) 0.741 0.852 0.703 0.862
150 Hz encodec [defossez2022high] 0.704 0.925 0.910 0.922
100 Hz dac [kumar2023high] 0.734 0.799 0.735 0.801
75 Hz audiodec [wu2023audiodec] 0.695 0.999 0.969 0.997
75 Hz flowdec [welker2025flowdec] 0.717 0.899 0.810 0.908
@@ Appendix C Foundation of bridge models
Bridge models provide a principled and efficient solution for inverse problems and conditional generation in settings where the prior is non-Gaussian yet partially aligned with the target distribution. Given a sample from the prior distribution $x_{\text{prior}}$ and its correspondence from target distribution $x_{\text{target}}$, In score-based generative models (SGMs) [song2020score], a forward SDE is defined between $x_{0}=x_{\text{target}}\sim p_{\text{target}}$, and $x_{T}=x_{\text{prior}}\sim p_{\text{prior}}$:
@@
Under the above framework with boundary constraints, the Schrödinger bridge (SB) problem [schrodinger1932theorie, chen2023schrodinger] seeks to find a path measure $p$ of specified boundary distributions that minimizes the Kullback-Leibler divergence $D_{\text{KL}}$ between a path measure and reference path measure $p_{\text{ref}}$:
@@
In SB theory [wang2021deep, chen2021likelihood, chen2023schrodinger], this specific SB problem can be expressed as a pair of forward-backwords linear SDEs:
@@
where the non-linear drifts $\nabla\log\Psi_{t}(x_{t})$ and $\nabla\log\widehat{\Psi}_{t}(x_{t})$ can be described by coupled partial differential equations (PDEs). A closed-form solution for SB [chen2023schrodinger] exists when Gaussian smoothing is applied with $p_{0}=\mathcal{N}\left(x_{\text{HR}},\epsilon_{0}^{2}I\right)$ and $p_{T}=\mathcal{N}\left(x_{\text{LR}},\epsilon_{T}^{2}I\right)$, to the original Dirac distribution. By defining $\alpha_{t}=e^{\int_{0}^{t}f(\tau)d\tau}$, $\bar{\alpha}_{t}=e^{\int_{1}^{t}f(\tau)d\tau}$, $\sigma_{t}^{2}=\int_{0}^{t}\frac{g^{2}(\tau)}{\alpha_{\tau}^{2}}d\tau$, and $\bar{\sigma}_{t}^{2}=\int_{t}^{1}\frac{g^{2}(\tau)}{\alpha_{\tau}^{2}}d\tau$, with boundary conditions and linear Gaussian assumption, tractable form of SB is solved as
@@
under $\epsilon_{T}=e^{\int_{0}^{T}f(\tau)d\tau}\epsilon_{0}$ and $\epsilon_{0}\rightarrow 0$ [chen2023schrodinger]. Therefore, marginal distribution of $x_{t}$ at state $t$ is also Gaussian:
@@ D.1 Waveform-based filtering augmentation
Figure 10: Visual comparison of waveform-based filtering strategies. (a) shows the baseline output from the any-to-48 kHz stage with high-frequency artifacts. (b) demonstrates that direct super-resolution from such inputs results in degraded high-frequency reconstruction. (c)–(e) apply low-pass filters with different cutoff frequencies, and the corresponding outputs after super-resolution are shown in (f)–(h).
Samples generated from the any-to-48 kHz stage often exhibit small-scale high-frequency artifacts, as shown in Figure 10 (a). Directly feeding such artifacts into the next-stage super-resolution model can significantly degrade performance, as shown in Figure 10 (b), where high-frequency content becomes difficult to reconstruct. To address this, we apply Chebyshev Type-I low-pass filters (order=8) with sharp frequency roll-off in the spectral domain, as shown in Figure 10 (c)–(e), removing a frequency content of 8 kHz, 4 kHz, and 2 kHz, respectively, before feeding them into the next-stage model. The resulting outputs are shown in Figure 10 (f)–(h).
@@ E.2 Low-frequency replacement
Low-frequency replacement is a widely adopted post-processing technique in super-resolution systems, particularly in super-resolution models based on compressed representations [liu2022neural, liu2024audiosr, yun2025flowhigh, im2025flashsr].
@@ E.3 Inference computational cost
Bridge-SR [li2025bridge] 48 kHz Waveform Bridge 1.670 50
UDM+ [liu2022neural] 48 kHz Waveform Unconditional Diffusion 2.320 100
AudioSR [liu2024audiosr] 48 kHz Mel-VAE Conditional Diffusion 0.948 50
Fre-painter [kim2024audio] 48 kHz Mel GAN 0.009 1
NVSR [liu2022neural] 48 kHz Mel GAN 0.033 1
@@
It is worth noting that under the 48 kHz setting, although our method is not as fast as GAN-based approaches such as [liu2022neural, kim2024audio], due to the iterative sampling nature, it significantly outperforms other iterative-based methods including [li2025bridge] and the previous state-of-the-art system AudioSR [liu2024audiosr].
When scaling to higher sampling rates (e.g., 192 kHz), it naturally leads to slower inference speeds. We mitigate this impact by adopting a lighter network architecture (see Appendix G.2), so the inference speed does not increase linearly with the sampling rate. As shown in Table 12, even when upsampling to 192 kHz, our system is still faster than the 48 kHz waveform-domain bridge system [li2025bridge].
@@ E.4 Modeling space and other probabilistic methods
To isolate the contribution of the latent-bridge formulation, we conducted additional experiments using the same speech datasets (OpenSLR [kjartansson2018crowd], Expresso [nguyen2023expresso], EARS [richter2023speech], and VCTK-Train [yamagishi2019cstr]) for training all models, and evaluated them on the VCTK-Test set. We compare the following variants without any additional modules, each trained with 500K steps on the any-to-48 kHz setting and tested on the 8-to-48 kHz SR setting with 50-step sampling:
@@
  • Bridge-STFT (based on the Nemo [Harper_NeMo_a_toolkit] framework)
  • Bridge-Waveform (based on Bridge-SR [li2025bridge])
@@
Method Modeling Space SSIM $\uparrow$ LSD $\downarrow$ LSD-LF $\downarrow$ LSD-HF $\downarrow$ SigMOS [ristea2025icassp] $\uparrow$
@@ G.1 Training dataset
48 kHz VCTK-train [yamagishi2019cstr] 40h Speech
OpenSLR [kjartansson2018crowd] 190h Speech
EARS [richter2023speech] 100h Speech
Expresso [nguyen2023expresso] 20h Speech
MusDB18 [rafii2017musdb18] 10h Music
Medleydb [bittner2014medleydb] 10h Music
@@ G.2 Model architecture
We adopt the Diffusion Transformer (DiT) architecture as the noise predictor for our bridge model, following the design of Stable Audio Open [evans2025stable]. During training, we employ the Bridge-gmax schedule, which has been proven effective in previous super-resolution and text-to-speech tasks [li2025bridge, chen2023schrodinger]. The network architecture and training configurations for the three-stage cascade are detailed below.
@@ G.3 Evaluation metrics
While LSD is a point-wise spectral metric, Structural Similarity (SSIM) [wang2004image] overcomes its limitations by incorporating structural and contextual information. Originally designed for perceptual image quality assessment, SSIM compares local statistics of luminance, contrast, and structure. In the context of audio, we apply SSIM on log-magnitude spectrograms to capture spectral texture fidelity.
@@
Here, $S_{k}$ and $\hat{S}_{k}$ denote the $k$-th local $7\times 7$ block in the spectrograms, $K$ is the total number of blocks, and $\epsilon_{1}=0.01$, $\epsilon_{2}=0.02$ are stability constants following VoiceFixer [du2024cosyvoice]. Higher SSIM values imply better preservation of spectral structure.
@@
To assess perceptual speech quality in an objective manner, we adopt the Virtual Speech Quality Objective Listener (ViSQOL) [chinen2020visqol], a signal-based estimator of Mean Opinion Score (MOS) using spectro-temporal similarity.
ViSQOL (in audio mode) assumes a sampling rate of 48 kHz and outputs MOS-LQO scores ranging from 1 to 4.75, with higher scores indicating better perceptual quality. However, since ViSQOL is designed for speech and only supports certain sample rates (e.g., 16 kHz), we additionally employ SigMOS [ristea2025icassp] to evaluate speech quality for 48 kHz scenario.
@@ G.4 Subjective experiment
For the subjective listening test, we evaluate super-resolution performance on generative model outputs from three representative systems: MaskGCT [wang2024maskgct], AudioLDM2 [liu2024audioldm], and QA-MDT [li2024quality], which represent modern speech generation at 24 kHz and music/sound generation at 16 kHz, respectively.
For speech, we randomly select 10 samples from the LibriSpeech-test set [panayotov2015librispeech] as both the reference audio and the text prompt for MaskGCT generation. For music and environmental audio, we sample 10 examples each from the MusicCaps [agostinelli2023musiclm] and AudioCaps [kim2019audiocaps] datasets. To further assess performance on real-world music, we additionally include 10 music segments from the Song-Describer dataset [manco2023song]. All audio clips are truncated to 5.12 seconds. For each test case, we present three versions to listeners: the input audio (Generated), the output from AudioSR, and the output from our system (AudioLBM), yielding a total of 40 groups × 3 samples = 120 audio clips for evaluation.
@@ H.1 Detailed introduction of audio super-resolution baselines
Nu-Wave [lee2021nu] represents the first attempt at applying diffusion models to waveform-level audio super-resolution. It builds on the WaveNet [van2016wavenet] structure and adopts the DiffWave [kong2020diffwave] training paradigm to learn mappings from fixed low-resolution audio to 48 kHz waveform. To support evaluation across various input resolutions, we follow the official implementation [^3] https://github.com/maum-ai/nuwave and train separate models for 8 kHz, 12 kHz, 16 kHz, and 24 kHz inputs, each for 1.5 million steps.
@@
Nu-Wave 2 [lee2021nu] extends Nu-Wave by incorporating two key innovations: Short-Time Fourier Convolution (STFC) and Bandwidth Spectral Feature Transform (BSFT). These additions improve harmonic modeling and allow support for any-to-48 kHz super-resolution. The model also adopts a fast generation scheme using eight predefined non-uniform sampling steps. For fair comparison, we use the pre-trained model made available by the authors [^4] https://github.com/maum-ai/nuwave2.
@@
UDM+ [yu2023conditioning] retains the DiffWave-based architecture but deviates from Nu-Wave in training strategy. It uses unconditional diffusion modeling and injects low-resolution signal guidance at inference through a 50-step uniform sampling process with low-frequency component replacement[lugmayr2022repaint]. Additionally, Manifold Constraint Gradient (MCG) [chung2022improving] is introduced to better balance low- and high-frequency consistency during generation.
@@
Bridge-SR [li2025bridge] adapts the Nu-Wave 2 architecture but introduces a key conceptual shift: instead of sampling from a standard Gaussian prior, the model initializes directly from the low-resolution input. This bridge formulation allows high-quality any-to-48 kHz SR. Furthermore, it employs data normalization strategies and frequency-aware loss functions to improve the final reconstruction quality.
@@
mdctGAN [shuai2023mdctgan] targets the instability issues of complex-valued neural networks in the STFT domain by operating in the Modified Discrete Cosine Transform (MDCT) domain. Rather than relying on post-vocoders, mdctGAN uses adversarial learning to generate high-fidelity waveform with phase consistency. We evaluate the model using the authors’ official checkpoints [^5] https://github.com/neoncloud/mdctGAN, covering four different input resolutions.
@@
NVSR [liu2022neural] is a neural vocoder-based speech SR system tailored to handle a wide range of upsampling factors and phase reconstruction challenges. It is composed of three main modules: (1) a mel-bandwidth extension network based on ResUNet; (2) a TFGAN-powered [tian2020tfgan] neural vocoder; and (3) a post-processing module. We use the official model checkpoint available at  [^6] https://github.com/haoheliu/ssr_eval for our comparisons.
@@
AP-BWE [lu2024towards] is the first bandwidth extension system to model the high-frequency phase explicitly. It leverages a GAN-based framework to jointly predict amplitude and phase spectra using a dual-stream CNN, where the two branches interact to reconstruct full-band audio from narrowband speech. Although it does not support arbitrary-resolution inputs, we evaluate the four fixed-resolution models provided by the authors [^7] https://github.com/yxlu-0102/AP-BWE.
@@
AudioSR [liu2024audiosr] is a general-purpose diffusion-based super-resolution model capable of handling speech, music, and general audio inputs. It performs super-resolution in the latent space conditioned on low-resolution inputs and uses a two-stage cascade—consisting of a mel-spectrogram VAE and a vocoder—for waveform reconstruction. Benefiting from strong zero-shot generalization, we evaluate the model using the official implementation [^8] https://github.com/haoheliu/versatile_audio_super_resolution.
@@
Fre-Painter [kim2024audio] introduces a masked autoencoder (MAE)-based framework for audio super-resolution. It employs an upper-band masking strategy during fine-tuning to simulate low-resolution inputs by masking high-frequency components. Additionally, a mix-ratio masking approach is utilized to enhance robustness across various input sampling rates. The model is pre-trained on large-scale datasets to learn robust speech representations and is fine-tuned jointly with a neural vocoder for waveform reconstruction. We evaluate the model using the official implementation [^9] https://github.com/FrePainter/code.
@@
FLowHigh [yun2025flowhigh] proposes an audio super-resolution method based on flow matching, aiming to overcome the sampling inefficiency of traditional diffusion models. It adopts an inpainting-style framework similar to AudioSR but constrains the modeling space to Mel-spectrograms. A lightweight two-layer Transformer [vaswani2017attention] is used as the velocity predictor. To further improve sampling efficiency, FLowHigh introduces a data-dependent prior instead of a pure Gaussian initialization, allowing for faster flow-based sampling. We evaluate the model using the official implementation [^10] https://github.com/jjunak-yun/FLowHigh_code.